OpenStudy (anonymous):

CALCULUS QUESTION! Temperature within a greenhouse over 24-hour period is given as T(t)=21+sin(pi t/12) a) determine the instantaneous rate of change at t=8... I found this by determining the derivative, and found the rate to be 0.85 b) determine the average rate of change of temp, within the greenhouse b/w the hours of 4-12. c) determine the avg temperature of greenhouse between the hours of 4-12. d) determine the average daily temp of the greenhouse.

OpenStudy (anonymous):

but the thing is i have to use integration.. but i don't understand how i would apply that concept to these questions.

OpenStudy (ipwnbunnies):

b. Avg. rate of change is found by using the slope formula. Finding the different between the values of the two points, and dividing that by the difference in the x-value. c. Average value of a closed interval can be calculated like this, if given a function: \[\frac{\int\limits_{a}^{b} f(x) dx}{b-a}\]

OpenStudy (anonymous):

uhm for d, since it's daily so 0-24 can't i just the same avg formula from b

OpenStudy (anonymous):

use**

OpenStudy (ipwnbunnies):

OH, it says 24 hour period lol. I believe you can still use that formula.

OpenStudy (ipwnbunnies):

Forget what I said about D earlier lol

OpenStudy (anonymous):

Alright! When I do average rate of change, do I use the original formula?

OpenStudy (ipwnbunnies):

What do you mean original formula? Average rate of change on a closed interval of a function is found by using the slope formula.

OpenStudy (anonymous):

I mean do I use T(t)=21+sin(pit/12) to find the corresponding temperature to the values t=4, and t=12 and then plug that into the slope formula?

OpenStudy (ipwnbunnies):

Yesh

OpenStudy (anonymous):

Okay! When i did that for d, i got 0. lol..

OpenStudy (anonymous):

which i don't think is right for the average daily temperature..

OpenStudy (ipwnbunnies):

I had a feeling it was 0, since there is the sine function there. It's a cyclical thing. When t = 0, you get sin(0) which is 0. When you plug in t = 24, sin(24pi/12) = sin(2pi) = 0

OpenStudy (ipwnbunnies):

That didn't really prove why it's 0, but even if you plug in t = 12, the value is 0. So, t = 0 to 24 is one cycle of a sine wave, and the average value of that is 0.

OpenStudy (anonymous):

Hmmm, alright. Thanks.

OpenStudy (anonymous):

:D